### v1

# Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback

Abstract
Feedback cooling plays a critical role in stabilizing quantum systems and achieving low temperatures, where a key question is to determine the fundamental thermodynamic limits on cooling performance. We establish a fundamental bound on quantum feedback cooling in Gaussian systems, by deriving a generalized second law of thermodynamics involving the kinetic temperatures of the system and a measure of quantum information flow obtained by continuous measurement. In contrast to previously known bounds, the obtained bound can be saturated by experimentally feasible situations using the quantum Kalman filter with a large feedback gain, where the cooling efficiency approaches its maximum. Our theoretical result is numerically demonstrated using parameters from an experiment of levitated nanoparticles. Our theory provides a general framework for understanding the thermodynamic constraints on quantum feedback cooling.
Introduction.— Measurement and feedback are fundamental tools that enable noise suppression, stabilization, optimal control, and adaptive control in diverse applications. In quantum systems, continuous measurement and feedback based on real-time data have been theoretically studied Wiseman and Milburn (2009); Jacobs and Steck (2006); Wiseman and Milburn (1993); Wiseman (1994); Doherty and Jacobs (1999); Bushev et al. (2006) and experimentally realized for qubit stabilization Vijay et al. (2012); Ristè et al. (2013), error correction Livingston et al. (2022), and cooling of oscillatory systems Rossi et al. (2018); Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023). In nanomechanical resonators Rossi et al. (2018) and optically trapped nanoparticles Gieseler et al. (2012); Magrini et al. (2021); Kamba et al. (2023); Gonzalez-Ballestero et al. (2021), state-based feedback control reduces vibrational energy, enabling cooling close to the ground state. In Gaussian systems, the quantum Kalman filter Belavkin (1995) provides optimal filtering that enhances the precision of feedback and has been demonstrated experimentally Magrini et al. (2021).
In feedback cooling, the energy of the system is reduced by extracting work through measurement-based control, which can be interpreted as a “Maxwell’s demon” setup in the context of thermodynamics of information Parrondo et al. (2015). By incorporating information terms, fundamental thermodynamic relations, such as the second law and the fluctuation theorem, have been generalized to systems under measurement and feedback in both classical and quantum regimes Sagawa and Ueda (2008); Jacobs (2009); Sagawa and Ueda (2010); Horowitz and Vaikuntanathan (2010); Funo et al. (2013); Gong et al. (2016), including the case of continuous measurement and feedback Sagawa and Ueda (2012); Ito and Sagawa (2013); Horowitz and Esposito (2014); Shiraishi and Sagawa (2015); Horowitz (2015); Ptaszyński and Esposito (2019); Yada et al. (2022). Experimental studies have further validated those information-thermodynamic relations Toyabe et al. (2010); Bérut et al. (2012); Koski et al. (2014); Ribezzi-Crivellari and Ritort (2019); Debiossac et al. (2020, 2022); Camati et al. (2016); Cottet et al. (2017); Masuyama et al. (2018); Naghiloo et al. (2018); Yada et al. (2024).
A fundamental challenge in feedback cooling is to determine the ultimate limits imposed by thermodynamics. While the standard formulation of the second law relates entropy production to the temperature of the heat bath, the primary objective in feedback cooling is to reduce the temperature of the system below the bath temperature. Thermodynamics of information for classical systems have revealed second-law-like inequalities for Langevin systems, which relate the kinetic temperatures of the system itself to the rate of information acquisition through measurement, quantified by transfer entropy Horowitz and Sandberg (2014); Sandberg et al. (2014); Hartich et al. (2014). However, a quantum counterpart to these results remains elusive, and a general characterization of cooling limits based on thermodynamic laws is still lacking for quantum systems.
In this Letter, we establish a fundamental bound of feedback cooling in Gaussian quantum systems. Specifically, we derive a second-law-like inequality that imposes a constraint on work extraction by a measure of quantum information flow called quantum-classical-transfer (QC-transfer) entropy Yada et al. (2022). Since the obtained bound is based on the kinetic temperatures of the system, it enables characterization of feedback cooling (i.e., reduction of the temperature of the system) and is achievable with a sufficiently large feedback gain along with using the quantum Kalman filter, in contrast to previously known bounds Yada et al. (2022). We numerically demonstrate these results with experimentally relevant parameters adopted from a recent nanoparticle experiment Magrini et al. (2021). Our results provide a theoretical foundation for the limits of feedback cooling from the perspective of thermodynamics of information in the quantum regime.
Figure 1: Schematic of the setup. The system undergoes feedback control based on estimated values obtained from the results of QND measurements. The system is also coupled to a heat bath at temperature $T$.
Setup.—We describe the framework for feedback cooling in Gaussian systems, where the state remains Gaussian throughout its evolution. We consider a quantum non-demolition (QND) continuous measurement and feedback control of a Gaussian system, which is also coupled to a single heat bath that satisfies the detailed balance condition (see Fig. 1). Specifically, the Hamiltonian of the system is supposed to be harmonic: $\hat{H}_{\mathrm{HO}}=\frac{\hbar\omega}{2}\left(\hat{p}^{2}+\hat{x}^{2}\right)$. Here, $\hat{x}$ and $\hat{p}$ represent the dimensionless position and momentum operators, respectively, satisfying the commutation relation $[\hat{x},\hat{p}]=i$, and $\omega$ is the angular frequency of the oscillator. We assume that the expectation values of the initial position and momentum are zero. The dynamics of feedback cooling are governed by the stochastic master equation Magrini et al. (2021):
$d\hat{\rho}_{\mathrm{c}}(t)=$ $-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{\mathrm{c}}(t)\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{\mathrm{c}}(t)]dt$
$+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}(t)]dt+\sqrt{2k\eta}\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}(t)]dW,$ (1)
where $\mathcal{D}_{\hat{z}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{z}\hat{\rho}_{\mathrm{c}}\hat{z}^{\dagger}-\frac{1}{2}\{\hat{z}^{\dagger}\hat{z},\hat{\rho}_{\mathrm{c}}\}$ is the dissipator and $\hat{\rho}_{\mathrm{c}}(t)$ represents the conditional state at time $t$ given previous measurement outcomes. The superoperator $\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\gamma(\bar{n}+1)\mathcal{D}_{\hat{a}}[\hat{\rho}_{\mathrm{c}}]+\gamma\bar{n}\mathcal{D}_{\hat{a}^{\dagger}}[\hat{\rho}_{\mathrm{c}}]$ represents the contribution of the heat bath, where $\hat{a}$ and $\hat{a}^{\dagger}$ represent the annihilation and creation operators, and $\gamma$ is the dissipation rate. The ratio between $\bar{n}$ and $\bar{n}+1$ satisfies the detailed balance condition $\bar{n}/(\bar{n}+1)=e^{-\hbar\omega/T}$, where $T$ is the temperature of the heat bath and the Boltzmann constant is set equal to unity for the sake of simplicity. The third term on the right-hand side of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) represents measurement backaction, while the fourth term reflects the reduction in position variance resulting from the measurement process. The superoperator $\mathcal{H}_{x}$ is defined as $\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]\coloneqq\hat{x}\hat{\rho}_{\mathrm{c}}+\hat{\rho}_{\mathrm{c}}\hat{x}-2\mathrm{tr}(\hat{x}\hat{\rho}_{\mathrm{c}})\hat{\rho}_{\mathrm{c}}$. The parameter $k$ represents the measurement rate, while the Wiener increment $dW$ is a Gaussian random variable with zero mean and variance $dt$. The measurement result obtained via QND measurement of the position over a short time interval $dt$ is described by $dy(t)=\braket{\hat{x}(t)}_{\mathrm{c}}dt+dW/\sqrt{8k\eta}$, where $\eta$ represents the detection efficiency and $\braket{\cdot(t)}_{\rm{c}}\coloneqq\mathrm{tr}\left(\hat{\rho}_{\rm{c}}(t)\;\cdot\right)$ denotes the expectation value under the conditional state. The set of all measurement results until time $t$ is defined as $Y_{t}\coloneqq\{dy(s)\}_{s=0}^{t^{-}}$. Note that the unconditional state is given by $\hat{\rho}=\mathbb{E}_{Y_{t}}[\hat{\rho}_{\rm c}]$, where $\mathbb{E}_{X}$ represents taking the expectation value over a random variable $X$.
Estimates of the observable are constructed from the set of obtained measurement outcomes and used for feedback. Using measurement results $Y_{t}$, estimates are defined as $e_{t}=h(Y_{t})$, where $h(\cdot)$ compresses the past measurement data into estimates $e_{t}$. The posterior state conditioned on this estimate is denoted by $\hat{\rho}_{t}^{e_{t}}$. Using the estimates, the feedback action is determined. The linear feedback Hamiltonian $\hat{H}_{\rm{FB}}$ is given by
$\hat{H}_{\mathrm{FB}}(t)=\hbar c_{x}(e_{t})\hat{x}+\hbar c_{p}(e_{t})\hat{p},$ (2)
where $c_{x}$ and $c_{p}$ are linear functions of the estimate $e_{t}$, defining the feedback coefficients. The special case of the estimate is the quantum Kalman filter Belavkin (1995), which is given by the conditional expectation values
$e_{t}=(\braket{\hat{x}(t)}_{\mathrm{c}},\braket{\hat{p}(t)}_{\mathrm{c}}).$ (3)
QC-transfer entropy.— As our goal is to obtain an information-thermodynamic relation to characterize the cooling limit of Gaussian systems, we consider the QC-transfer entropy introduced in Ref. Yada et al. (2022), which quantifies the amount of information gain through continuous measurement. It is defined as
$\dot{I}_{\mathrm{QCT}}\coloneqq\lim_{dt\rightarrow 0}\frac{1}{dt}\mathbb{E}_{Y_{t}}\left[I_{\mathrm{QC}}\left(\hat{\rho}_{\mathrm{c}}:dy(t)\right)\right].$ (4)
Here, $I_{\mathrm{QC}}\left(\hat{\rho}_{\mathrm{c}}:dy(t)\right)$ is the quantum-classical-mutual (QC-mutual) information Groenewold (1971); Ozawa (1986) given by
$I_{\mathrm{QC}}\left(\hat{\rho}_{\mathrm{c}}:dy(t)\right)\coloneqq S\left(\hat{\rho}_{\mathrm{c}}\right)-\mathbb{E}_{dy(t)}\left[S\left(\hat{\rho}_{\mathrm{c,\mathrm{M}}}\right)\right],$ (5)
where $\hat{\rho}_{\mathrm{c,\mathrm{M}}}:=\hat{\rho}_{\mathrm{c}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}]dt+4k\eta\mathcal{H}_{\hat{x}}[\hat{\rho}_{\mathrm{c}}](dy(t)-\braket{\hat{x}(t)}_{\mathrm{c}}dt)$ is the conditional state after measurement at time $t$, and $S(\hat{\rho})\coloneqq-\mathrm{tr}[\hat{\rho}\ln\hat{\rho}]$ represents the von Neumann entropy. We note that Eq. (5) is regarded as the conditional QC-mutual information given all prior measurement results $Y_{t}$, because $\hat{\rho}_{\rm c}$ is the conditional state. Since the QC-transfer entropy quantifies the total information gain, it provides an upper bound for the information available in the multiple feedback processes.
Figure 2: Schematic of the distribution of a Gaussian state $\hat{\rho}(t)$ in the phase space (i.e., the Wigner distribution). The kinetic temperatures $T_{u}$ and $T_{v}$ characterize the variances in the eigen-directions of the covariance matrix $\bm{\sigma}$.
[…]
$T_{u}\coloneqq{\rm{e}}^{2r}T^{\mathrm{G}},\quad T_{v}\coloneqq{\rm{e}}^{-2r}T^{\mathrm{G}}.$ (6)
A schematic of kinetic temperatures is shown in Fig. 2. The kinetic temperatures $T_{u}$ and $T_{v}$ correspond to scaling the Gibbs state temperature $T^{\mathrm{G}}$ according to the widths in eigen-directions of the distribution.
We next define the work rate as follows. First, the internal energy of the system is given by $E(t)=\mathrm{tr}[\hat{H}_{\mathrm{HO}}\hat{\rho}(t)]$. Then, based on the first law of thermodynamics, the work rate corresponding to the energy extracted from the system is defined as $\dot{W}_{\mathrm{ext}}:=\dot{Q}-\frac{d}{dt}E(t)$, where $\dot{Q}:=\mathrm{tr}\left(\hat{H}_{\mathrm{HO}}\mathcal{L}_{\mathrm{bath}}[\rho]\right)$ is the heat Gong et al. (2016). The time variation of the energy can be expressed as contributions along the directions of $\sigma_{uu}$ and $\sigma_{vv}$, using the quadrature operators $\hat{u}$ and $\hat{v}$ associated with these directions, as $\frac{d}{dt}E(t)=\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{u}^{2}\right]+\frac{\hbar\omega}{2}\frac{d}{dt}\mathrm{tr}\left[\hat{\rho}\hat{v}^{2}\right]$. Similarly, the work rate can be further decomposed as:
$\dot{W}_{\mathrm{ext}}=\dot{W}_{\mathrm{ext},u}+\dot{W}_{\mathrm{ext},v},$ (7)
where the terms $\dot{W}_{\mathrm{ext},u}$ and $\dot{W}_{\mathrm{ext},v}$ represent the contributions of the work rate along the eigen-directions of $\bm{\sigma}$.
Main results.— We now present the main result of this study, a fundamental bound of the work rate in terms of kinetic temperatures. Suppose that the system is in the steady state; the unconditional state $\hat{\rho}$ is given by the steady solution of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). Then, we have a generalized second law of thermodynamics in the form
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}\leq\dot{I}_{\mathrm{QCT}}.$ (8)
This inequality indicates how much QC-transfer entropy can be utilized to enhance work extraction and lower the kinetic temperatures, and serves as a quantum generalization of the second-law-like bound for classical Langevin systems Horowitz and Sandberg (2014).
We compare our result (8) with the generalized second law using the temperature $T$ of the heat bath, given in the form $\dot{W}_{\rm{ext}}/T\leq\dot{I}_{\mathrm{QCT}}$ Yada et al. (2022). When feedback cooling reduces the kinetic temperatures $T_{u}$ and $T_{v}$ below the bath temperature $T$, we have $\dot{W}_{\rm{ext}}/T\leq\dot{W}_{\mathrm{ext},u}/T_{u}+\dot{W}_{\mathrm{ext},v}/T_{v}$. In such a case, inequality (8) provides a tighter bound, as it explicitly involves the kinetic temperatures of the system itself. Since cooling is achieved by extracting work from the system and lowering its kinetic temperatures, inequality (8) represents the fundamental cooling limit.
We next consider the conditions for achieving the equality in (8). A key ingredient is the quantum Kalman filter, which provides the conditional expectation values as the estimates, as presented in Eq. (3). Note that $\hat{\rho}_{t}^{e_{t}}=\hat{\rho}_{\mathrm{c}}$ holds in this case. By using this estimate based on the quantum Kalman filter, one can implement the so-called state-based feedback Doherty and Jacobs (1999); Magrini et al. (2021), where the feedback Hamiltonian (2) is given by the coefficients $c_{x}=b_{x}\langle\hat{x}\rangle_{\mathrm{c}}+b_{p}\langle\hat{p}\rangle_{\mathrm{c}}$ and $c_{p}=-a_{x}\langle\hat{x}\rangle_{\mathrm{c}}-a_{p}\langle\hat{p}\rangle_{\mathrm{c}}$. This feedback Hamiltonian aims to reduce the mean position and momentum to the origin in the phase space. Another requirement for achieving the equality in (10) is a large feedback gain. In fact, the equality in (10) is asymptotically achieved as the feedback gain goes to infinity if the estimate is implemented by the quantum Kalman filter (see the Supplemental Material sup for the proof). For simplicity, let the parameters of the feedback Hamiltonian be $a_{x}=b_{p}=g$ and $a_{p}=b_{x}=0$. In this case, the equality in (10) is asymptotically achieved in the limit $g\to\infty$. This implies that, under these conditions, the feedback system perfectly converts the acquired information into work.
Under the same conditions, the occupancy number of the harmonic Hamiltonian $\hat{H}_{\rm HO}$, $\langle a^{\dagger}a\rangle$, also reaches its minimal value determined by the bath temperature and the measurement rate (see Supplemental Material sup ), as the optimal feedback minimizes the sum of the variances of position and momentum, which is equivalent to the occupation number in our setup. This highlights the connection between the information-theoretic bound on thermodynamic energy conversion and the cooling limit of Gaussian systems.
Outline of the proof, and more on quantum information flow.— To outline the derivation of our main inequality (8) and to clarify the structure behind its equality condition, here we introduce yet another information quantity, named the quantum-classical (QC) information flow, which is slightly different from the QC-transfer entropy. The QC information flow is defined as
$\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}\coloneqq\lim_{dt\rightarrow 0}\frac{1}{dt}\biggl{[}\chi(\hat{\rho}_{t+dt}:e_{t})-\chi(\hat{\rho}_{t}:e_{t})\biggr{]},$ (9)
where $\chi(\rho:e)\coloneqq S(\rho)-\mathbb{E}_{e}\left[S(\rho^{e})\right]$ is the Holevo information Holevo (1973). Since the Holevo information is the measure of the correlation between a quantum state and a classical memory, the QC information flow $-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}$ quantifies the change in the system-memory correlation induced by the time evolution of the system during the feedback. We note that the QC information flow is a quantum version of the classical information flow Horowitz and Esposito (2014), which is defined as the partial derivative of the classical mutual information.
Based on the QC information flow, we can derive a generalized second law that is also applicable to non-steady states:
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}\leq-\dot{I}_{\mathrm{QCI}}^{S}-\dot{S}_{\mathrm{BA}}.$ (10)
Here, $\dot{S}_{\mathrm{BA}}$ represents the entropy change due to measurement backaction, defined as
$\dot{S}_{\mathrm{BA}}:=\lim_{dt\rightarrow 0}\frac{1}{dt}\mathbb{E}_{e_{t}}\left[\left(S\left(\hat{\rho}_{t,\mathrm{BA}}^{e_{t}}\right)-S\left(\hat{\rho}_{t}^{e_{t}}\right)\right)\right],$ (11)
where $\hat{\rho}^{e_{t}}_{t,\mathrm{BA}}:=\hat{\rho}_{t}^{e_{t}}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{t}^{e_{t}}]dt$ is the state evolved for a time interval $dt$ under the contributions of measurement backaction. The term $-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}-\dot{S}_{\mathrm{BA}}$ quantifies the information available for feedback control, because the unavoidable entropy change due to the measurement backaction is subtracted from the change in correlation originating from the dynamics of the system. The detailed derivation of (10) is provided in the Supplemental Material sup . A sufficient condition to achieve the equality of (10) is the quantum Kalman filter and a large feedback gain.
We next relate the two measures of quantum information flow: QC-transfer entropy and QC information flow. Again suppose that the system is in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). Then, the QC-transfer entropy provides an upper bound of the QC information flow:
$-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}-\dot{S}_{\mathrm{BA}}\leq\dot{I}_{\mathrm{QCT}}.$ (12)
See Supplemental Material sup for the derivation. A sufficient condition for the equality in (12) is the quantum Kalman filter. For the classical case of (12), the equality condition can be expressed in terms of sufficient statistics Matsumoto and Sagawa (2018).
Now, by combining inequalities (10) and (12), our main result (8) is obtained. The equality condition for (8) directly follows from the equality conditions for inequalities (10) and (12).
Figure 3: The behavior of ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ as a function of the feedback gain $g$ in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). The feedback protocol is based on the quantum Kalman filter, where the feedback Hamiltonian is given by $H^{\hat{x},\hat{p}}_{\rm{FB}}=\hbar g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$ (orange curve). In addition, we demonstrate another feedback Hamiltonian $H^{\hat{x}}_{\rm{FB}}=\hbar g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$ (blue dashed curve) for the sake of comparison. The behavior of ${\dot{W}_{\rm{ext}}}/{T}$ is also shown (green curve). The black line represents the QC-transfer entropy, the bound of  (8), to which only the orange curve approaches in large $g$. The parameters are chosen as $k/\omega=0.18$, $\eta=0.34$, $T=292$K, and $\bar{n}\gamma/\omega=0.0058$ Magrini et al. (2021). The vertical axis is normalized by $\dot{I}_{\mathrm{QCT}}$
Numerical demonstration.— We present a numerical demonstration of our main inequality (8), by numerically implementing the quantum Kalman filter. Figure 3 shows the behavior of ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ and ${\dot{W}_{\rm{ext}}}/{T}$ as functions of the feedback gain in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). The parameters are taken from Ref. Magrini et al. (2021), an experiment of feedback cooling of levitated nanoparticles.
Figure 3 compares the two feedback Hamiltonians and shows that only $H^{\hat{x},\hat{p}}_{\rm{FB}}=\hbar g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$, plotted in the orange curve, ensures that ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ approaches the upper bound $\dot{I}_{\mathrm{QCT}}$, as $g\to\infty$. Since feedback proportional to $\hat{x}$ is commonly used in experiments, we also consider the Hamiltonian of the form $H^{\hat{x}}_{\rm{FB}}=\hbar g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$ (blue dashed curve), but it does not allow the system to reach the information-theoretic bound of (8). Moreover, $\dot{W}_{\rm{ext}}/T$ with feedback Hamiltonian $H^{\hat{x},\hat{p}}_{\rm{FB}}$ (green curve) fails to asymptotically approach $\dot{I}_{\rm QCT}$, indicating that the naive bound without using the kinetic temperatures is not achievable.
Conclusion and outlook.— In this study, we established a fundamental limit of quantum feedback cooling in Gaussian systems under continuous QND measurement, by linking work extraction, kinetic temperatures, and QC-transfer entropy (8). Since the derived inequality is based on the kinetic temperatures rather than the bath temperature, it imposes a stricter constraint on cooling performance than known information-thermodynamic inequalities, and can be saturated through a sufficiently large feedback gain with the quantum Kalman filter (Fig. 3). In addition, we discussed the relationship between QC-transfer entropy and yet another quantity called QC information flow, as characterized by inequality (12). These results elucidate to what extent information acquired from the measurement can enhance the extraction of the work and reduce the kinetic temperatures.
Our results can be experimentally verified using current or near-future experimental technologies. For example, in levitated nanoparticle systems, feedback cooling based on the quantum Kalman filter has been realized Magrini et al. (2021), making experimental verification promising. However, current experiments typically utilize feedback proportional only to the position operator, whereas achieving the theoretical limit requires feedback proportional to the momentum operator as well. Realizing this optimal feedback is experimentally challenging, but is anticipated to lead to the verification of the fundamental connection between information and thermodynamics in feedback cooling.
Acknowledgements.
Acknowledgments.— We are grateful to Kiyotaka Aikawa and Isaac Layton for valuable discussions. This work was supported by JST ERATO Grant No. JPMJER2302, Japan. K.K. and T.Y. are supported by World-leading Innovative Graduate Study Program for Materials Research, Information, and Technology (MERIT-WINGS) of the University of Tokyo. T.Y. is also supported by JSPS KAKENHI Grant No. JP23KJ0672. K.F. is supported by JSPS KAKENHI Grant Nos. JP23K13036 and JP24H00831. T.S. is supported by JSPS KAKENHI Grant No. JP19H05796 and JST CREST Grant No. JPMJCR20C1. T.S. is also supported by Institute of AI and Beyond of the University of Tokyo.
Supplemental Material for “Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback”
Kousuke Kumasaki1, Toshihiro Yada1, Ken Funo1, and Takahiro Sagawa1,2
1Department of Applied Physics, The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-8656, Japan
2Quantum-Phase Electronics Center (QPEC), The University of Tokyo, 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-8656, Japan
I Quantum information flow in general setup
In this section, we introduce the measures of quantum information flow and derive the generalized second law, in the setup of multiple quantum measurement and feedback, which includes the continuous feedback case, mainly addressed in this manuscript, as a special case. In Sec. I.1, we introduce QC information flow, and derive the generalized second law. In Sec. I.2, we introduce QC-transfer entropy and discuss its relationship with QC information flow. In Sec. I.3, we demonstrate that the continuous feedback control setup discussed in the main text is a special case of the general multiple feedback setup.
I.1 Generalized second law with QC information flow
Figure S1: Schematic for the multiple measurement and feedback setup. We consider the situation where the measurement and feedback are alternately performed. The $n$-th measurement outcome is represented as $y_{t_{n}}$, and all the measurement outcomes until $y_{t_{n}}$ is denoted as $Y_{t_{n}}\coloneqq(y_{t_{0}},y_{t_{1}},\dots,y_{t_{n}})$. The state right before and after the $n$-th measurement are denoted as $\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$ and $\hat{\rho}_{t_{n+1}}^{Y_{t_{n+1}}}$, respectively, whose superscript represents the corresponding measurement outcomes. The superoperator for $n$-th feedback process is described as $\mathcal{E}_{t_{n}}^{e_{t_{n}}}$, where $e_{t_{n}}$ represents the estimator determined by $Y_{t_{n}}$.
We here consider the general setup of multiple measurement and feedback, as shown in Fig. S1. The process starts at the time $t=t_{0}$, and the $n$-th measurement and feedback is conducted during the time interval $[t_{n},t_{n+1})$. The $n$-th measurement outcome is represented as $y_{t_{n}}$, and all the measurement outcomes until $y_{t_{n}}$ is denoted as $Y_{t_{n}}\coloneqq(y_{t_{0}},y_{t_{1}},\dots,y_{t_{n}})$. The conditional density operators right before and after the $n$-th measurement are denoted as $\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$ and $\hat{\rho}_{t_{n+1}}^{Y_{t_{n+1}}}$, respectively, whose superscript represents the corresponding measurement outcomes. The $n$-th feedback process depends on the $n$-th estimator, which is determined as a function of $Y_{t_{n}}$ as $e_{t_{n}}\coloneqq h(Y_{t_{n}})$, and $\mathcal{E}_{t_{n}}^{e_{t_{n}}}$ is the superoperator representing the feedback process. The density operator conditioned on the estimator $e_{t_{n}}$ is represented as $\hat{\rho}_{t_{n}}^{e_{t_{n}}}$, where the relationship $\sum_{h(Y_{t_{n}})=e_{t_{n}}}P(Y_{t_{n}})\hat{\rho}_{t_{n}}^{Y_{t_{n}}}=P(e_{t_{n}})\hat{\rho}_{t_{n}}^{e_{t_{n}}}$ is satisfied. The feedback process consists of the heat-bath interaction and external driving, and therefore the superoperator can be decomposed as $\mathcal{E}_{t_{n}}^{e_{t_{n}}}=\mathcal{E}^{e_{t_{n}}}_{\rm drv}\circ\mathcal{E}^{e_{t_{n}}}_{\rm bath}$.
In such a setup, we derive the generalized second law during for the time interval spanning from the $(n_{i}+1)$-th to the $n_{f}$-th measurement and feedback process. The quantum-classical (QC) information flow during this period can be described as
$I_{\mathrm{QCI}}^{S}\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\chi\left(\hat{\rho}_{t_{n+1}}:e_{t_{n}}\right)-\chi\left(\hat{\rho}_{t_{n}}:e_{t_{n}}\right),$ (S1)
where $\chi(\hat{\rho}:e)\coloneqq S(\hat{\rho})-\mathbb{E}_{e}[S(\hat{\rho}^{e})]$ represents the Holevo information Holevo (1973), $S(\hat{\rho})\coloneqq-\mathrm{tr}[\hat{\rho}\ln\hat{\rho}]$ is the von Neumann entropy, $\mathbb{E}_{X}$ represents taking the expectation value over a random variable $X$, and $\hat{\rho}_{t_{n}}\coloneqq\mathbb{E}_{e_{t_{n}}}[\hat{\rho}_{t_{n}}^{e_{t_{n}}}]$. Since the Holevo information quantifies the correlation between the quantum system and the (classical) estimator, $-I_{\mathrm{QCI}}^{S}$ is the total information utilized in the multiple feedback. The effect of measurement backaction is quantified as
$S_{\rm BA}\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}\left[S\left(\hat{\rho}_{t_{n+1}}^{e_{t_{n}}}\right)-S\left(\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)\right].$ (S2)
Since $\hat{\rho}_{t_{n+1}}^{e_{t_{n}}}$ is described as $\hat{\rho}_{t_{n+1}}^{e_{t_{n}}}=\mathcal{M}_{\mathrm{BA},t_{n}}(\hat{\tau}_{t_{n}}^{e_{t_{n}}})$, where $\mathcal{M}_{\mathrm{BA},t_{n}}$ is the superoperator representing the measurement backaction, $S_{\rm BA}$ is the accumulation of the entropy increase due to the measurement backaction.
By denoting the system Hamiltonian as $H$, the averaged heat transferred from the heat bath is defined as
$Q\coloneqq\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}[Q^{e_{t_{n}}}]=\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{e_{t_{n}}}\left[\mathrm{tr}\left[H(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})-\hat{\rho}^{e_{t_{n}}}_{t_{n}})\right]\right],$ (S3)
which quantifies the total energy increase due to the interaction with the heat bath. The extracted work and free energy change can also be defined as $W_{\rm ext}\coloneqq-E(\hat{\rho}_{t_{n_{f}}})+E(\hat{\rho}_{t_{n_{i}}})+Q,$ and $\Delta F\coloneqq E(\hat{\rho}_{t_{n_{f}}})-E(\hat{\rho}_{t_{n_{i}}})-\beta^{-1}(S(\hat{\rho}_{t_{n_{f}}})-S(\hat{\rho}_{t_{n_{i}}})),$ respectively, where $E(\hat{\rho})\coloneqq\mathrm{tr}[H\hat{\rho}]$ represents the internal energy of the system.
By explicitly describing the expectation value of the function $g$, $\mathbb{E}_{e}[g(e_{t_{n}})]$, as the sum $\sum_{e_{t_{n}}}P[e_{t_{n}}]g(e_{t_{n}})$, we can derive the generalized second law as follows:
$\begin{split}\beta(W_{\rm ext}+\Delta F)+I_{\mathrm{QCI}}^{S}+S_{\rm BA}&=-S(\hat{\rho}_{t_{n_{f}}})+S(\hat{\rho}_{t_{n_{i}}})+\beta Q+I_{\mathrm{QCI}}^{S}+S_{\rm BA}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S\left(\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)-S\left(\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-\beta Q^{e_{t_{n}}}\right]\right\}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S\left(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})\right)-S\left(\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-\mathrm{tr}\left[\left(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})-\hat{\rho}^{e_{t_{n}}}_{t_{n}}\right)\ln\hat{\rho}^{\mathrm{G}}\right]\right]\right\}\\ &=-\sum_{n={n_{i}}}^{n_{f}-1}\left\{\mathbb{E}_{e_{t_{n}}}\left[S(\hat{\rho}^{e_{t_{n}}}_{t_{n}}\|\hat{\rho}^{\mathrm{G}})-S(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})\|\hat{\rho}^{\mathrm{G}})\right]\right\}\\ &\leq 0,\end{split}$ (S4)
where $S(\hat{\rho}\|\hat{\sigma})\coloneqq\mathrm{tr}[\hat{\rho}(\ln\hat{\rho}-\ln\hat{\sigma})]$ represents the quantum relative entropy, and $\hat{\rho}^{\mathrm{G}}\propto\exp(-\beta H)$ is a Gibbs state with the inverse temperature $\beta$. The equality in the third line follows from the fact that $\hat{\tau}_{t_{n}}^{e_{t_{n}}}=\mathcal{E}^{e_{t_{n}}}_{\rm drv}\circ\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}})$ with the superoperator $\mathcal{E}^{e_{t_{n}}}_{\rm drv}$ representing purely unitary driving, and therefore $S(\hat{\tau}_{t_{n}}^{e_{t_{n}}})=S(\mathcal{E}^{e_{t_{n}}}_{\rm bath}(\hat{\rho}^{e_{t_{n}}}_{t_{n}}))$. The inequality in the final line follows from the fact that $\hat{\rho}^{\mathrm{G}}$ is the fixed point of the heat-bath interaction $\mathcal{E}^{e_{t_{n}}}_{\rm bath}$, and the monotonicity of quantum relative entropy under a CPTP (completely positive and trace-preserving) map Petz (2003).
I.2 QC-transfer entropy
The QC-transfer entropy Yada et al. (2022) for the time interval from the $(n_{i}+1)$-th to $n_{f}$-th process is described as
$I_{\mathrm{QCT}}\coloneqq\sum_{n={n_{i}}}^{n_{f}-1}\mathbb{E}_{Y_{t_{n}}}\left[I_{\mathrm{QC}}\left(\hat{\tau}_{t_{n}}^{Y_{t_{n}}}:y_{t_{n+1}}\right)\right],$ (S5)
where $I_{\mathrm{QC}}$ represents the quantum-classical mutual (QC-mutual) information Groenewold (1971); Ozawa (1986), defined as $I_{\mathrm{QC}}\left(\hat{\tau}:y\right)\coloneqq S(\hat{\tau})-\mathbb{E}_{y}[S(\hat{\rho}^{y})]$, with the post-selected state $\hat{\rho}^{y}$ for measurement outcome $y$. The QC-mutual information $I_{\mathrm{QC}}$ quantifies the information obtained by a single measurement, and the QC-transfer entropy $I_{\mathrm{QCT}}$ is the total information transferred through the multiple measurements.
The QC-transfer entropy and the QC information flow satisfies the inequality
$I_{\mathrm{QCT}}-\Delta\chi\geq-I_{\mathrm{QCI}}^{\mathrm{S}}-S_{\mathrm{BA}},$ (S6)
where $\Delta\chi\coloneqq\chi(\hat{\rho}_{t_{n_{f}}}:Y_{t_{n_{f}}})-\chi(\hat{\rho}_{t_{n_{i}}}:Y_{t_{n_{i}}})$ is the change in correlation from $n=n_{i}$ to $n=n_{f}$. By explicitly describing the expectation value $\mathbb{E}_{Y_{t_{n}}}[g(Y_{t_{n}})]$ as the sum $\sum_{Y_{t_{n}}}P[Y_{t_{n}}]g(Y_{t_{n}})$, we can derive this inequality as follows:
$\begin{split}I_{\mathrm{QCT}}-\Delta\chi+I_{\mathrm{QCI}}^{\mathrm{S}}+S_{\rm BA}&=\sum_{n=n_{i}}^{n_{f}-1}\mathbb{E}_{Y_{t_{n}}}\left[S\left(\hat{\rho}_{t_{n}}^{Y_{t_{n}}}\|\hat{\rho}_{t_{n}}^{e_{t_{n}}}\right)-S\left(\hat{\tau}_{t_{n}}^{Y_{t_{n}}}\|\hat{\tau}_{t_{n}}^{e_{t_{n}}}\right)\right]\\ &\geq 0,\end{split}$ (S7)
where the first line follows from a straightforward rearrangement of terms, and the second line follows from the equalities $\mathcal{E}_{t_{n}}^{e_{t_{n}}}(\hat{\rho}_{t_{n}}^{e_{t_{n}}})=\hat{\tau}_{t_{n}}^{e_{t_{n}}}$ and $\mathcal{E}_{t_{n}}^{e_{t_{n}}}(\hat{\rho}_{t_{n}}^{Y_{t_{n}}})=\hat{\tau}_{t_{n}}^{Y_{t_{n}}}$, along with the monotonicity of quantum relative entropy under a CPTP map. As can be seen from Eq. (S7), a simple equality condition for Eq. (S6) is that $\hat{\rho}_{t_{n}}^{e_{t_{n}}}=\hat{\rho}_{t_{n}}^{Y_{t_{n}}}$ holds for all $Y_{t_{n}}$. This condition implies that the feedback estimator $e_{t_{n}}=h(Y_{t_{n}})$ is optimally designed, which is known to be achieved by a Kalman filter.
I.3 Correspondence to the continuous feedback setup considered in the main text
We note that the definitions of thermodynamic quantities and information terms in the main text are consistent with those in Secs. I.1 and I.2. In particular, the consistency of the heat income $Q$, work extraction $W_{\rm ext}$, and QC information flow $I_{\mathrm{QCI}}^{\rm S}$ can be directly verified by substituting the discrete time index $t_{n}$ with the continuous time variable $t$ as is demonstrated in Sec. II.2. While the correspondence is less straightforward for the measurement backaction term $S_{\mathrm{BA}}$ [Eq. (11) in the main text] and the QC-transfer entropy $I_{\mathrm{QCT}}$ [Eq. (4) in the main text], it can still be demonstrated by neglecting terms of order $O(dt)$. It is comprehensive to represent Eqs. (S2) and (S5) in the following forms taking the limit of $dt\to 0$:
$\dot{S}_{\mathrm{BA}}$ $=\mathbb{E}_{e_{t}}\left[\left.\lim_{dt\to 0}\frac{S(\mathcal{M}_{\mathrm{BA},t}(\hat{\rho}))-S(\hat{\rho})}{dt}\right|_{\hat{\rho}=\hat{\tau}_{t}^{e_{t}}}\right],$ (S8)
$\dot{I}_{\mathrm{QCT}}$ $=\mathbb{E}_{Y_{t}}\left[\left.\lim_{dt\to 0}\frac{I_{\rm QC}(\hat{\rho}:dy(t))}{dt}\right|_{\hat{\rho}=\hat{\tau}_{t}^{Y_{t}}}\right].$ (S9)
From the equalities $\hat{\tau}_{t}^{e_{t}}=\hat{\rho}_{t}^{e_{t}}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{e_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{e_{t}}]dt$ and $\hat{\tau}_{t}^{Y_{t}}=\hat{\rho}_{t}^{Y_{t}}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{Y_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{Y_{t}}]dt,$ and the fact that the terms in the square brackets are smooth at the respective points $\hat{\rho}=\hat{\tau}_{t}^{e_{t}}$ and $\hat{\rho}=\hat{\tau}_{t}^{Y_{t}}$, we can replace $\hat{\tau}_{t}^{e_{t}}$ and $\hat{\tau}_{t}^{Y_{t}}$ in these equations with $\hat{\rho}_{t}^{e_{t}}$ and $\hat{\rho}_{t}^{Y_{t}}$, respectively, introducing only negligible additional terms of order $O(dt)$. We also note that in the main text, we use the notations $\hat{\rho}_{\rm c}(t)=\hat{\rho}_{t}^{Y_{t}}$ and $\hat{\rho}_{t,\mathrm{BA}}^{e_{t}}=\mathcal{M}_{\mathrm{BA},t}(\hat{\rho}_{t}^{e_{t}})$ for simplicity. In Sec. II, we use the expressions (S8) and (S9) to derive the main results presented in the main text.
II Information thermodynamic cooling limit in Gaussian systems
In this section, we present the derivation of inequality (10) in the main text:
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}\leq-\dot{I}_{\mathrm{QCI}}^{S}-\dot{S}_{\rm BA}.$ (S10)
We further show the following relation (12) in the main text under the steady-state condition:
$-\dot{I}_{\mathrm{QCI}}^{S}-\dot{S}_{\rm BA}\leq\dot{I}_{\rm QCT}.$ (S11)
Inequalities (10) and (S11) are derived for a Gaussian system that obeys the following stochastic master equation:
$d\hat{\rho}_{t}^{Y_{t}}=-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}}+\hat{H}_{\mathrm{FB}}(t),\hat{\rho}_{t}^{Y_{t}}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}_{t}^{Y_{t}}]dt+2k\mathcal{D}_{\hat{x}}[\hat{\rho}_{t}^{Y_{t}}]dt+\sqrt{2k\eta}\mathcal{H}_{\hat{x}}[\hat{\rho}_{t}^{Y_{t}}]dW.$ (S12)
We note that the dynamics of the system described by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) is a special case of the general setup discussed in Sec. I.1, where the explicit form of the superoperators for the heat-bath interaction, external driving, and measurement backaction are respectively given by:
$\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho})$ $=\hat{\rho}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{HO}},\hat{\rho}\right]dt+\mathcal{L}_{\mathrm{bath}}[\hat{\rho}]dt,$ (S13)
$\mathcal{E}^{e_{t}}_{\rm drv}(\hat{\rho})$ $=\hat{\rho}-\frac{i}{\hbar}\left[\hat{H}_{\mathrm{FB}}(t),\hat{\rho}\right]dt,$ (S14)
$\mathcal{M}_{\mathrm{BA},t}(\hat{\rho})$ $=\hat{\rho}+2k\mathcal{D}_{\hat{x}}[\hat{\rho}]dt,$ (S15)
and the time interval $dt\coloneqq t_{n+1}-t_{n}$ is taken to the limit $dt\to 0$. Using these superoperators, the time-evolution of the quantum state $\rho_{t}^{e_{t}}$ from $t$ to $t+dt$ for given $e_{t}$ reads:
$\hat{\rho}_{t}^{e_{t}}\rightarrow\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}_{t}^{e_{t}})=:\hat{\pi}_{t}^{e_{t}}\rightarrow\mathcal{E}^{e_{t}}_{\rm drv}(\hat{\pi}_{t}^{e_{t}})=:\hat{\tau}_{t}^{e_{t}}\rightarrow\mathcal{M}_{\mathrm{BA},t}(\hat{\tau}_{t}^{e_{t}}):=\hat{\rho}_{t+dt}^{e_{t}}.$ (S16)
We note that the Hamiltonian of the system $H$ is given by $\hat{H}_{\rm HO}=(\hbar\omega/2)(\hat{x}^{2}+\hat{p}^{2})$ and the feedback Hamiltonian is defined as $\hat{H}_{\mathrm{FB}}(t)=\hbar c_{x}(e_{t})\hat{x}-\hbar c_{p}(e_{t})\hat{p}$, where $c_{x}$ and $c_{p}$ are linear functions of estimate $e_{t}$, defining the feedback coefficients. We further assume that the feedback Hamiltonian is chosen so that, when the initial state satisfies $\langle\hat{x}\rangle(0)=\langle\hat{p}\rangle(0)=0$, the expectation values at later times satisfy $\langle\hat{x}\rangle(t)=\langle\hat{p}\rangle(t)=0$. Intuitively, such condition is satisfied when the feedback Hamiltonian is designed to reduce the mean position and momentum to the origin of the phase space. On the other hand, such condition is not satisfied when the feedback Hamiltonian realizes a constant displacement of the position $\hat{H}_{\mathrm{FB}}(t)=a\hat{x}$, where $a$ is a constant.
In what follows, we first derive the representation of kinetic temperatures using the Gaussian property of the quantum state in Sec. II.1, and then derive (10) in Sec. II.2. We further assume the steady-state condition and derive (S11) in Sec. II.3.
II.1 Kinetic temperatures
In what follows, we introduce the kinetic temperatures of one-mode Gaussian state $T_{u}$ and $T_{v}$. We start by noting that any single-mode Gaussian states centered at the origin can be regarded as a squeezed thermal state Kim et al. (1989) obtained by applying a squeezing operation $\hat{S}(\xi)$ to the Gibbs state $\hat{\rho}^{\mathrm{G}}$. In other words, for any single-mode Gaussian state $\hat{\rho}_{t}$ centered at the origin, there exists a pair $(\hat{\rho}^{\mathrm{G}},\xi)$ such that Olivares (2012):
$\hat{\rho}_{t}=\hat{S}(\xi)\hat{\rho}^{\mathrm{G}}\hat{S}(\xi)^{\dagger}.$ (S17)
Here, the squeezing operator is parameterized by the complex number $\xi=r\mathrm{e}^{i\theta}$ as
$\hat{S}(\xi):=\exp\left(\frac{1}{2}\xi(\hat{a}^{\dagger})^{2}-\frac{1}{2}\xi^{*}\hat{a}^{2}\right),$ (S18)
and the Gibbs state is expressed in terms of the inverse temperature $\beta^{\mathrm{G}}$ as
$\hat{\rho}^{\mathrm{G}}:=\mathrm{e}^{-\beta^{\mathrm{G}}\hat{H}_{\mathrm{HO}}}/Z(\beta^{\mathrm{G}}),$ (S19)
where $Z(\beta)$ is the partition function given by $Z(\beta^{\mathrm{G}}):=\mathrm{tr}[\mathrm{e}^{-\beta^{\mathrm{G}}\hat{H}_{\mathrm{HO}}}]$. Now, the explicit expression of the covariance matrix of the Gaussian state reads
$\sigma_{xx}$ $=[\cosh{(2r)}+\sinh{(2r)}\cos\theta]\sigma^{\rm G},$
$\sigma_{pp}$ $=[\cosh{(2r)}-\sinh{(2r)}\cos\theta]\sigma^{\rm G},$
$\sigma_{xp}$ $=\sinh{(2r)}\sin\theta\sigma^{\rm G},$ (S20)
where
$\sigma^{\mathrm{G}}:=\mathrm{tr}(\hat{x}^{2}\hat{\rho}^{\rm G})=\mathrm{tr}(\hat{p}^{2}\hat{\rho}^{\rm G})=\frac{1}{2}\coth\frac{\hbar\omega\beta^{\rm G}}{2}$ (S21)
is the common eigenvalues of the covariance matrix of the Gibbs state. Then, the eigenvalues of the covariance matrix of $\hat{\rho}_{t}$ satisfy
$\sigma_{uu}=\mathrm{e}^{2r}\sigma^{\mathrm{G}},\quad\sigma_{vv}=\mathrm{e}^{-2r}\sigma^{\mathrm{G}}.$ (S22)
Thus, each eigenvalue of the covariance matrix of $\hat{\rho}$ corresponds to the variance of the Gibbs state scaled by $\mathrm{e}^{2r}$ or $\mathrm{e}^{-2r}$ using the squeeze parameter.
We shall now define the kinetic temperatures of one-mode Gaussian state assosiated with the eigenvalues $\sigma_{uu}$ and $\sigma_{vv}$ as
$T_{u}\coloneqq\mathrm{e}^{2r}T^{\mathrm{G}},\quad T_{v}\coloneqq\mathrm{e}^{-2r}T^{\mathrm{G}},$ (S23)
where
$T^{\mathrm{G}}=\hbar\omega\left[\ln\left(\frac{2\sqrt{\det\bm{\sigma}}+1}{2\sqrt{\det\bm{\sigma}}-1}\right)\right]^{-1}.$ (S24)
We further note that
$\ln\hat{\rho}_{t}$ $=\hat{S}(\xi)\left(\ln\hat{\rho}^{\mathrm{G}}\right)\hat{S}(\xi)^{\dagger}$
$=-\beta^{\rm G}\hat{S}(\xi)\hat{H}_{\rm HO}\hat{S}(\xi)^{\dagger}-\ln Z(\beta^{\mathrm{G}})$
$=-\frac{\hbar\omega\hat{u}^{2}}{2T_{u}}-\frac{\hbar\omega\hat{v}^{2}}{2T_{v}}-\ln Z(\beta^{\mathrm{G}}),$ (S25)
where
$\hat{u}$ $\coloneqq e^{r}\hat{S}(\xi)\hat{x}\hat{S}(\xi)^{\dagger},$ (S26)
$\hat{v}$ $\coloneqq e^{-r}\hat{S}(\xi)\hat{p}\hat{S}(\xi)^{\dagger},$ (S27)
are the quadratures that correspond to the eigenvalues $\sigma_{uu}$ and $\sigma_{vv}$, i.e., $\sigma_{uu}=\mathrm{tr}(\hat{u}^{2}\hat{\rho}_{t})-\mathrm{tr}(\hat{u}\hat{\rho}_{t})^{2}$ and $\sigma_{vv}=\mathrm{tr}(\hat{v}^{2}\hat{\rho}_{t})-\mathrm{tr}(\hat{v}\hat{\rho}_{t})^{2}$, respectively.
II.2 Derivation of the information thermodynamic cooling limit
In the following, we derive (10). We start by noting that from the definition of the heat (S3), the heat rate reads
$\dot{Q}:=\mathbb{E}_{e_{t}}\left[\lim_{dt\rightarrow 0}\frac{1}{dt}\Bigl{\{}\mathrm{tr}[(\hat{\pi}_{t}^{e_{t}}-\hat{\rho}_{t}^{e_{t}})\hat{H}_{\mathrm{HO}}]\Bigr{\}}\right].$ (S28)
Similarly, the work rate reads:
$\dot{W}_{\mathrm{ext}}$ $:=-\mathbb{E}_{e_{t}}\left[\lim_{dt\rightarrow 0}\frac{1}{dt}\Bigl{\{}\mathrm{tr}\left[(\hat{\rho}_{t+dt}^{e_{t}}-\hat{\pi}_{t}^{e_{t}})\hat{H}_{\mathrm{HO}}\right]\Bigr{\}}\right].$ (S29)
We then define the work rate along $\hat{u}$ (S26) and $\hat{v}$ (S27) as:
$\dot{W}_{\mathrm{ext},u}$ $:=-\frac{\hbar\omega}{2}\mathbb{E}_{e_{t}}\left[\lim_{dt\rightarrow 0}\frac{1}{dt}\Bigl{\{}\mathrm{tr}\left[(\hat{\rho}_{t+dt}^{e_{t}}-\hat{\pi}_{t}^{e_{t}})\hat{u}^{2}\right]\Bigr{\}}\right],$ (S30)
$\dot{W}_{\mathrm{ext},v}$ $:=-\frac{\hbar\omega}{2}\mathbb{E}_{e_{t}}\left[\lim_{dt\rightarrow 0}\frac{d}{dt}\Bigl{\{}\mathrm{tr}\left[(\hat{\rho}_{t+dt}^{e_{t}}-\hat{\pi}_{t}^{e_{t}})\hat{v}^{2}\right]\Bigr{\}}\right].$ (S31)
We now combine with the expression (II.1) and find that the left-hand side of (10) is expressed as
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}$ $=\mathbb{E}_{e_{t}}\left[\lim_{dt\rightarrow 0}\frac{d}{dt}\Bigl{\{}\mathrm{tr}\left[(\hat{\rho}_{t+dt}^{e_{t}}-\hat{\pi}_{t}^{e_{t}})\ln\hat{\rho}_{t}\right]\Bigr{\}}\right]$
$=\lim_{dt\rightarrow 0}\frac{d}{dt}\Bigl{[}\mathrm{tr}\left[(\hat{\rho}_{t+dt}-\hat{\pi}_{t})\ln\hat{\rho}_{t}\right]\Bigr{]}$
$=\lim_{dt\rightarrow 0}\frac{d}{dt}\Bigl{[}-S(\hat{\rho}_{t+dt})+S(\hat{\pi}_{t})\Bigr{]},$ (S32)
where we used the fact that $S(\hat{\rho}_{t+dt})=-\mathrm{tr}[\hat{\rho}_{t+dt}\ln\hat{\rho}_{t}]+O(dt^{2})$, etc.
We next discuss the QC information flow (S1) for the time evolution given by (S16), which reads
$\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}$ $:=\lim_{dt\rightarrow 0}\frac{\chi\left(\hat{\rho}_{t+dt}:e_{t}\right)-\chi\left(\hat{\rho}_{t}:e_{t}\right)}{dt}.$ (S33)
We further decompose Eq. (S33) into each step in (S16), as
$\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}=\dot{I}_{\mathrm{QCI,B}}^{\mathrm{S}}+\dot{I}_{\mathrm{QCI,D}}^{\mathrm{S}}+\dot{I}_{\mathrm{QCI,BA}}^{\mathrm{S}},$ (S34)
where each term is defined as:
$\dot{I}_{\mathrm{QCI,B}}^{\mathrm{S}}$ $:=\lim_{dt\rightarrow 0}\frac{\chi\left(\hat{\pi}_{t}:e_{t}\right)-\chi\left(\hat{\rho}_{t}:e_{t}\right)}{dt},$ (S35)
$\dot{I}_{\mathrm{QCI,D}}^{\mathrm{S}}$ $:=\lim_{dt\rightarrow 0}\frac{\chi\left(\hat{\tau}_{t}:e_{t}\right)-\chi\left(\hat{\pi}_{t}:e_{t}\right)}{dt},$ (S36)
$\dot{I}_{\mathrm{QCI,BA}}^{\mathrm{S}}$ $:=\lim_{dt\rightarrow 0}\frac{\chi\left(\hat{\rho}_{t+dt}:e_{t}\right)-\chi\left(\hat{\tau}_{t}:e_{t}\right)}{dt}.$ (S37)
Since the von Neumann entropy of the system is unchanged under the unitary time evolution (S14), i.e., $S(\hat{\pi}_{t}^{e_{t}})=S(\hat{\tau}_{t}^{e_{t}})$, Eq. (S36) simplifies to
$\dot{I}_{\mathrm{QCI,D}}^{\mathrm{S}}=\lim_{dt\rightarrow 0}\frac{S(\hat{\tau}_{t})-S(\hat{\pi}_{t})}{dt}.$ (S38)
From Eqs. (S34) and (S38), we obtain the relation:
$\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}=\dot{I}_{\mathrm{QCI,B}}^{\mathrm{S}}-\dot{S}_{\mathrm{BA}}+\lim_{dt\rightarrow 0}\frac{S\left(\hat{\rho}_{t+dt}\right)-S\left(\hat{\pi}_{t}\right)}{dt},$ (S39)
where the entropy change of the system due to the measurement backaction $\dot{S}_{\rm BA}$ is given by Eq. (S8). We now show the following relation:
$\dot{I}_{\mathrm{QCI,B}}^{\mathrm{S}}\leq 0,$ (S40)
which follows from
$\chi\left(\hat{\pi}_{t}:e_{t}\right)-\chi\left(\hat{\rho}_{t}:e_{t}\right)$ $=S(\mathcal{E}_{\rm bath}(\hat{\rho}_{t}))-S(\hat{\rho}_{t})-\mathbb{E}_{e_{t}}\Bigl{[}S(\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}_{t}^{e_{t}}))-S(\hat{\rho}_{t}^{e_{t}}))\Bigr{]}$
$=\mathbb{E}_{e_{t}}\Bigl{[}S(\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}_{t}^{e_{t}})\|\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}_{t}))-S(\hat{\rho}_{t}^{e_{t}}\|\hat{\rho}_{t})\Bigr{]}$
$\leq 0.$ (S41)
Here, we used the relation $\mathbb{E}_{e_{t}}[\mathcal{E}^{e_{t}}_{\rm bath}(\hat{\rho}^{e_{t}}_{t})]=\mathcal{E}^{e_{t}}_{\rm bath}(\mathbb{E}_{e_{t}}[\hat{\rho}_{t}^{e_{t}}])$ which directly follows from Eq. (S13), and the monotonicity of the quantum relative entropy and obtained the last inequality in (II.2). By combining Eqs. (S39), (II.2), and (II.2), we have
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}=-\lim_{dt\rightarrow 0}\frac{S\left(\hat{\rho}_{t+dt}\right)-S\left(\hat{\pi}_{t}\right)}{dt}\leq-\dot{I}_{\mathrm{QCI}}^{\mathrm{S}}-\dot{S}_{\rm BA},$ (S42)
which completes the proof of (10).
II.3 Derivation of the inequality relation between the QC information flow and the QC-transfer entropy
We shall now prove (S11) by considering an infinitesimal time-step shown in Eq. (S16) and use the relation (S7):
$\dot{I}_{\rm QCT}-\dot{\chi}+\dot{I}_{\rm QCI}^{\rm S}+\dot{S}_{\rm BA}\geq 0,$ (S43)
where the QC-transfer entropy $\dot{I}_{\rm QCT}$ is given by Eq. (S9), and the infinitesimal variation of the Holevo information reads
$\dot{\chi}:=\lim_{dt\rightarrow 0}\frac{1}{dt}\Bigl{[}S(\hat{\rho}_{t+dt})-\mathbb{E}_{Y_{t+dt}}[S(\hat{\rho}^{Y_{t+dt}}_{t+dt})]-S(\hat{\rho}_{t})+\mathbb{E}_{Y_{t}}[S(\hat{\rho}_{t}^{Y_{t}})]\Bigr{]}.$ (S44)
In what follows, we show that $\dot{\chi}$ vanishes in the steady-state of the Gaussian time-evolution described by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback), thereby proving (S11). In doing so, we first note that the von Neumann entropy of a single-mode Gaussian state can be expressed using its covariance matrix $\bm{\sigma}$ as Olivares (2012)
$S(\hat{\rho}_{t})=f\left(\sqrt{\det\bm{\sigma}}\right),$ (S45)
where $f$ is a single-variable function defined as
$f(x)\coloneqq\left(x+\frac{1}{2}\right)\ln\left(x+\frac{1}{2}\right)-\left(x-\frac{1}{2}\right)\ln\left(x-\frac{1}{2}\right).$ (S46)
In particular, the von Neumann entropy of the conditional state $\hat{\rho}_{t}^{Y_{t}}$ is solely determined by $\bm{\sigma}_{\rm c}(t)\coloneqq[\sigma_{ij,\mathrm{c}}(t)]_{i,j=x,p}$, where $\sigma_{xx,\mathrm{c}}(t)\coloneqq\mathrm{tr}(\hat{x}^{2}\hat{\rho}_{t}^{Y_{t}})-[\mathrm{tr}(\hat{x}\hat{\rho}_{t}^{Y_{t}})]^{2}$, and similar definition applies to $\sigma_{pp,\mathrm{c}}(t)$ and $\sigma_{xp,\mathrm{c}}(t)$. For Gaussian dynamics described by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback), we note that $\bm{\sigma}_{\rm c}$ does not depend on the measurement outcome $dy$, as their time-evolution explicitly reads:
$\frac{d}{dt}\sigma_{xx,{\mathrm{c}}}$ $=2\omega\sigma_{xp,{\mathrm{c}}}-\gamma\sigma_{xx,{\mathrm{c}}}+\frac{\gamma}{2}(2\bar{n}+1)-8k\eta\sigma_{xx,{\mathrm{c}}}^{2}$ (S47)
$\frac{d}{dt}\sigma_{pp,\mathrm{c}}$ $=-2\omega\sigma_{xp,{\mathrm{c}}}-\gamma\sigma_{pp,{\mathrm{c}}}+\frac{\gamma}{2}(2\bar{n}+1)+2k-8k\eta\sigma_{xp,{\mathrm{c}}}^{2}$ (S48)
$\frac{d}{dt}\sigma_{xp,{\mathrm{c}}}$ $=-\omega\sigma_{xx,{\mathrm{c}}}+\omega\sigma_{pp,{\mathrm{c}}}-\gamma\sigma_{xp,{\mathrm{c}}}-8k\eta\sigma_{xx,{\mathrm{c}}}\sigma_{xp,{\mathrm{c}}}.$ (S49)
Therefore, in the steady-state, we have
$\hat{\rho}_{t}=\hat{\rho}_{t+dt},\ \ \ \bm{\sigma}_{\rm c}(t)=\bm{\sigma}_{\rm c}(t+dt).$ (S50)
By combining Eqs. (S43), (S44) and (S50), we complete the proof of (S11). As discussed in Sec. I.2, the equality in (S11) is achieved by a quantum Kalman filter.
II.4 Technical notes
For a single-mode Gaussian state, the kinetic temperature introduced in this study plays the role of converting entropy into energy. For higher-dimensional systems or non-Gaussian states, it may be possible to generalize this relationship by extending the concept of kinetic temperature. However, a comprehensive method for defining kinetic temperatures in such scenarios remains elusive and requires further exploration.
III Cooling limit
In Sec. III.1, we discuss the cooling limit of the occupation number (S51) and the equality condition of the generalized second law (8) based on the quantum Kalman filter. In Sec. III.2, we consider a direct feedback scheme, in which the measurement outcome $dy$ itself is used as an estimate, and show numerical plots of Eqs. (S51) and (8).
III.1 Achievement of the cooling limit based on the quantum Kalman filter
An example of experimental systems that obey the stochastic master equation (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) is the levitated nanoparticle system Magrini et al. (2021). In nanoparticle systems, the degree of cooling is often evaluated by how small the occupation number of the particle, $\langle n\rangle=\langle a^{\dagger}a\rangle$, can be made. Using the variances of the state, this can be expressed as $\braket{n}=\frac{1}{2}\left(\sigma_{xx}+\sigma_{pp}\right)-1/2$ Defining the occupation number corresponding to $\bm{\sigma}_{\mathrm{c}}$ as $\langle n\rangle_{\min}\coloneqq\frac{1}{2}(\sigma_{xx,\mathrm{c}}+\sigma_{pp,\mathrm{c}})$, we obtain:
$\braket{n}\geq\braket{n}_{\mathrm{min}}:=\frac{1}{2}\left(\sigma_{xx,\mathrm{c}}^{2}+\sigma_{pp,\mathrm{c}}^{2}\right)-\frac{1}{2}.$ (S51)
Here, $\langle n\rangle_{\min}$ represents the cooling limit. The equation of the inequality (S51) is fulfilled when $\bm{\sigma}=\bm{\sigma}_{\mathrm{c}}$ holds. In the absence of a heat bath, the equality condition for (S51) is known to be satisfied in the limit of strong feedback using the quantum Kalman filter estimates Doherty and Jacobs (1999). A similar sufficient condition can also be demonstrated in this case. At the same time, the equality condition of the inequality (8) in the main text
$\frac{\dot{W}_{\mathrm{ext},u}}{T_{u}}+\frac{\dot{W}_{\mathrm{ext},v}}{T_{v}}\leq\dot{I}_{\mathrm{QCT}}$ (S52)
holds. This means that the information thermodynamic bound obtained in this study and the cooling limit of nanoparticle systems are reached simultaneously. In what follows, we show the equality conditions of Eqs. (S51) and (8).
We now briefly introduce the quantum Kalman filter, in which the estimated value $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$ is obtained. We assume that a feedback Hamiltonian $\hat{H}_{\mathrm{FB}}(t)=-\left(a_{x}\braket{x}_{\rm{c}}+a_{p}\braket{p}_{\rm{c}}\right)\hat{p}+\left(b_{x}\braket{x}_{\rm{c}}+b_{p}\braket{p}_{\rm{c}}\right)\hat{x}$ is applied using this estimated value. Then, time evolutions of the expected values $\braket{\hat{x}}_{\mathrm{c}}$ and $\braket{\hat{p}}_{\mathrm{c}}$ governed by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback) are expressed as:
$d\braket{\hat{x}}_{\mathrm{c}}=$ $\left(\braket{\hat{p}}_{\mathrm{c}}-a_{x}\braket{\hat{x}}_{\mathrm{c}}-a_{p}\braket{\hat{p}}_{\mathrm{c}}-\frac{\gamma}{2}\braket{\hat{x}}_{\mathrm{c}}\right)dt+8k\eta\sigma_{xx,\mathrm{c}}\cdot(dy-\braket{x}_{\mathrm{c}}dt),$ (S53)
$d\braket{\hat{p}}_{\mathrm{c}}=$ $\left(-\omega\braket{\hat{x}}_{\mathrm{c}}-b_{x}\braket{\hat{x}}_{\mathrm{c}}-b_{p}\braket{\hat{p}}_{\mathrm{c}}-\frac{\gamma}{2}\braket{\hat{p}}_{\mathrm{c}}\right)dt+8k\eta\sigma_{xp,\mathrm{c}}\cdot(dy-\braket{x}_{\mathrm{c}}dt).$ (S54)
This stochastic differential equation allows us to determine the expected values of the position and momentum at the next time step based on current expected values and the measurement outcome $dy$. In the quantum Kalman filter, this mechanism is used to sequentially obtain the estimate $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$. Here, the variance and covariance of the posterior state appearing in the filtering coefficients can be determined in advance because the differential equations of $\bm{\sigma}_{\mathrm{c}}$ are closed with respect to $\bm{\sigma}_{\mathrm{c}}$ [see Eq. (S49)].
We further analyze the fluctuations inherent in the estimates of the quantum Kalman filter $e_{t}=\{\braket{\hat{x}}_{\mathrm{c}},\braket{\hat{p}}_{\mathrm{c}}\}$. The variance and covariance of the estimates are defined as $\sigma_{xx,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}^{2}]-(\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}])^{2}$, $\sigma_{pp,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}^{2}]-(\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}])^{2}$, $\sigma_{xp,{\mathrm{m}}}\coloneqq\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}\braket{\hat{p}}_{\mathrm{c}}]-\mathbb{E}_{e_{t}}[\braket{\hat{x}}_{\mathrm{c}}]\mathbb{E}_{e_{t}}[\braket{\hat{p}}_{\mathrm{c}}]$. We define the covariance matrix of these as $\bm{\sigma}_{\mathrm{m}}$. The covariance matrix can be decomposed as $\bm{\sigma}=\bm{\sigma}_{\mathrm{c}}+\bm{\sigma}_{\mathrm{m}}$. From Eqs. (S53) and (S54), the differential equations governing the time evolution of $\bm{\sigma}_{\mathrm{m}}$ are obtained as:
$\frac{d}{dt}\sigma_{xx,{\mathrm{m}}}$ $=2\omega\sigma_{xp,{\mathrm{m}}}-2a_{x}\sigma_{xx,{\mathrm{m}}}-2a_{p}\sigma_{xp,{\mathrm{m}}}-\gamma\sigma_{xx,{\mathrm{m}}}+8k\eta\sigma_{xx,\mathrm{c}}^{2}$ (S55)
$\frac{d}{dt}\sigma_{pp,{\mathrm{m}}}$ $=-2\omega\sigma_{xp,{\mathrm{m}}}-2b_{x}\sigma_{xp,{\mathrm{m}}}-2b_{p}\sigma_{pp,{\mathrm{m}}}-\gamma\sigma_{pp,{\mathrm{m}}}+8k\eta\sigma_{xp,{\mathrm{c}}}^{2}$ (S56)
$\frac{d}{dt}\sigma_{xp,{\mathrm{m}}}$ $=-(\omega+b_{x})\sigma_{xx,{\mathrm{m}}}+(\omega-a_{p})\sigma_{pp,{\mathrm{m}}}-(a_{x}+b_{p}+\gamma)\sigma_{xp,{\mathrm{m}}}+8k\eta\sigma_{xx,\mathrm{c}}\sigma_{xp,\mathrm{c}}$ (S57)
In these equations, terms involving $a_{x}$, $a_{p}$, $b_{x}$, $b_{p}$ represent the contributions from feedback control. We note that the steady-state value of $\bm{\sigma}_{\rm c}$ does not depend on feedback gains (S49). Therefore, the reduction of $\braket{n}$ is solely determined by how much $\bm{\sigma}_{\mathrm{m}}$ can be reduced via feedback control. For simplicity, let $a_{x}=b_{p}=g$, $a_{p}=b_{x}=0$, and assume that $g$ is a constant. Under these conditions, it can be shown that, in the limit $g\to\infty$, each component of $\bm{\sigma}_{\mathrm{m}}$ scales as $O(1/g)$. This implies that $\bm{\sigma}_{\mathrm{m}}$ can be arbitrarily small in the limit of an infinite feedback rate. Since $\bm{\sigma}\rightarrow\bm{\sigma}_{\mathrm{c}}$ holds in this regime, the cooling limit determined by (S51) is achieved. In this case, $\hat{\rho}_{t}^{e_{t}}=\hat{\rho}_{t}$ holds, thus achieving the equality in (II.2), and also the equality in (10). Moreover, since Eq. (S11) holds in the steady state with quantum Kalman filter, the equality in (8) is also achieved.
We now prove the asymptotic property $\bm{\sigma}_{\mathrm{m}}=O(g)$. The time evolution equations for $\bm{\sigma}_{\mathrm{m}}$, as described by Eqs. (S55), (S56), and (S57), can be rewritten in vector form as follows:
$\frac{d}{dt}\bm{s}_{\mathrm{m}}=\bm{A}\bm{s}_{\mathrm{m}}+\bm{v},$ (S58)
where
$\bm{s}_{\mathrm{m}}:=[\sigma_{xx,{\mathrm{m}}},\sigma_{pp,{\mathrm{m}}},\sigma_{xp,{\mathrm{m}}}]^{\mathrm{T}}$ (S59)
represents the variances and covariances of $\bm{\sigma}_{\mathrm{m}}$. Here, $\bm{v}\coloneqq[8k\eta\sigma_{xx,\mathrm{c}}^{2},8k\eta\sigma_{xp,{\mathrm{c}}}^{2},8k\eta\sigma_{xx,\mathrm{c}}\sigma_{xp,\mathrm{c}}]^{\mathrm{T}}$, and the matrix $\bm{A}$ is given by
$\bm{A}\coloneqq\left(\begin{array}[]{ccc}-2a_{x}-\gamma&0&2\omega\\ 0&-2b_{p}-\gamma&-2\omega\\ -\omega&\omega&-a_{x}-b_{p}-\gamma\end{array}\right).$
The eigenvalues of the $\bm{A}$ are given by
$\lambda_{1}=-(a_{x}+b_{p}+\gamma),$ $\lambda_{2}=-(a_{x}+b_{p}+\gamma)+\sqrt{(a_{x}-b_{p})^{2}-4\omega^{2}},$ $\lambda_{3}=-(a_{x}+b_{p}+\gamma)-\sqrt{(a_{x}-b_{p})^{2}-4\omega^{2}}.$ (S60)
Using a unitary matrix $\bm{U}$, $\bm{A}$ can be diagonalized as $\bm{U}^{\dagger}\bm{A}\bm{U}=\mathrm{diag}(\lambda_{1},\lambda_{2},\lambda_{3})$. Defining $\tilde{\bm{s}}\coloneqq\bm{U}\bm{s}_{\mathrm{m}}$ and $\tilde{\bm{v}}\coloneqq\bm{U}\bm{v}$, the components of $\tilde{\bm{s}}$ satisfy:
$\frac{d}{dt}\tilde{s}_{i}(t)=\lambda_{i}\tilde{s}_{i}(t)+\tilde{v}_{i}(t).$ (S61)
Assuming $\tilde{s}_{i}(0)=0$, the solution is:
[…]
Applying the triangle inequality to Eq. (S61), we obtain
$\left|\tilde{s}_{i}(t)\right|\leq\frac{\left|\tilde{v}_{i,\max}\right|}{\left|\lambda_{i}\right|}\left|1-\mathrm{e}^{\lambda_{i}t}\right|,$ (S63)
where $|\tilde{v}_{i,\max}|=\max{0\leq\tau\leq t}|\tilde{v}_{i}(\tau)|$. For $a_{x}=b_{p}=g$, we have $1/\lambda_{i}=O(1/g)$, so the right-hand side of (S63) is $O(1/g)$. Therefore, $\tilde{\bm{s}}(t)=O(1/g)$. Since $\bm{s}_{\mathrm{m}}=\bm{U}^{\dagger}\tilde{\bm{s}}$, we conclude that $\bm{s}_{\mathrm{m}}=O(1/g)$.
Figure S2: The behavior of average occupation number $\braket{n}$ (left plot) and ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ (right plot) as a function of the feedback gain $g$ in the steady state of Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). The feedback protocol is based on the quantum Kalman filter, where the feedback Hamiltonian is given by $H^{\hat{x},\hat{p}}_{\rm{FB}}=\hbar g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$ (orange curve). In addition, we demonstrate another feedback Hamiltonian $H^{\hat{x}}_{\rm{FB}}=\hbar g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$ (blue dashed curve). In the left plot, the black line represents cooling limit $\braket{n}_{\mathrm{min}}$. In the right plot, the black line represents the QC-transfer entropy, and the vertical axis is normalized by $\dot{I}_{\mathrm{QCT}}$. The parameters are chosen as $k/\omega=0.18$, $\eta=0.34$, $T=292$K, and $\bar{n}\gamma/\omega=0.0058$  Magrini et al. (2021).
We present a numerical demonstration regarding the achievement of the cooling limit. Figure S2 illustrates the behavior of the average occupation number $\langle n\rangle$ as functions of the feedback gain in the steady state described by Eq. (Thermodynamic approach to quantum cooling limit of continuous Gaussian feedback). We compute the behavior for the two Hamiltonians $H^{\hat{x},\hat{p}}_{\rm{FB}}=g(-\langle x\rangle_{\rm{c}}\hat{p}+\langle p\rangle_{\rm{c}}\hat{x})$ and $H^{\hat{x}}_{\rm{FB}}=g(\langle x\rangle_{\rm{c}}+\langle p\rangle_{\rm{c}})\hat{x}$. It is evident that when using $H^{\hat{x},\hat{p}}_{\rm{FB}}$, the cooling limit $\langle n\rangle_{\rm{min}}$ is asymptotically achieved as the feedback gain $g$ approaches infinity. In contrast, for $H^{\hat{x}}_{\rm{FB}}$, the occupation number $\langle n\rangle$ asymptotically approaches a value larger than the cooling limit $\langle n\rangle_{\rm{min}}$. In practical experiments, feedback proportional to $\hat{x}$ is commonly employed. However, such feedback does not allow the system to reach the cooling limit $\langle n\rangle_{\rm{min}}$. A similar trend is observed in the relationship between ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ and QC-transfer entropy.
III.2 Direct Feedback
Finally, we consider the dynamics of direct feedback using the QND measurement result $dy$. In other words, $dy$ itself is used as an estimate in this feedback scheme. The time-evolution for the direct feedback is given by
$\mathcal{E}^{e_{t}}_{\rm drv}(\hat{\rho}_{\rm c}(t))=\hat{\rho}_{\rm c}(t)-\frac{i}{\hbar}[-a_{x}\hat{p}+b_{x}\hat{x},\hat{\rho}_{\rm c}(t)]\circ dy(t-\tau_{\rm FB}),$ (S64)
where $a_{x}$ and $b_{x}$ are the feedback gains, $\circ$ denote the Stratonovitch product, and $\tau_{\rm FB}$ is the feedback delay time. We consider the limit of $\tau_{\rm FB}\rightarrow 0$ that corresponds to the case of Markovian feedback Wiseman (1994); Wiseman and Milburn (2009).
Figure S3: Behavior of the steady-state average occupation number of the harmonic oscillator, $\braket{n}$ (left plot), and the quantity ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ (right plot) as functions of the feedback gain $g$ under direct feedback. The parameters are the same as in Fig. S2. Therefore, the values of $\braket{n}_{\rm min}$ and $\dot{I}_{\rm QCT}$ are identical to those in Fig. S2.
Figure S3 illustrates the behavior of the steady-state average occupation number $\braket{n}$ of the harmonic oscillator and the quantity ${\dot{W}_{\mathrm{ext},u}}/{T_{u}}+{\dot{W}_{\mathrm{ext},v}}/{T_{v}}$ as functions of the feedback gain $g$ under direct feedback. Unlike the case where the quantum Kalman filter is used, the equality in (S11) does not hold for direct feedback, and the cooling limit is not achieved. Additionally, it is observed that the best cooling is achieved at a finite value of the feedback gain $g$, in contrast to the case of using the quantum Kalman filter, where the optimal cooling is achieved when the feedback gain becomes infinite.